OpenStudy (cicilybailey14):

My grandson is about as many days as my son in weeks, and my grandson is as many months as I am in years. My grandson, my son and I together are 120 years. Can you tell me my age in years

OpenStudy (danjs):

You need to keep everything in the same units, since the answer is in years, we will let each variable be in terms of years... Let G=grandson age in years S=Son age in years I = Self age in years

OpenStudy (danjs):

Then, translate each given sentence into an equation.. My grandson is about as many days as my son in weeks, remember the variables are in years... days of granson 365*G is same as Son weeks 52*S 365*G = 52*S grandson is as many months as I am in years 12*G = I My grandson, my son and I together are 120 years. G + S + I = 120

OpenStudy (danjs):

you can solve those 3 equations as a system and get values for ages in years, 365*G = 52*S 12*G = I G + S + I = 120 ------------------- if you want a decimal approximation -- S=42.07 yrs , G=5.99 yrs, I=71.93 yrs